# A simple explanation for the absence of the spherical nuclei

### v1

@@ Abstract
Since the EMC effect increases with the number of nucleons, it can be guessed that the effect of this color degree of freedom is more obvious for heavy nuclei. The fact is true. The SU3-IBM explains the prolate-oblate asymmetric phase transition in the Hf-Hg region, where the SU(3) third-order Casimir operator describes the oblate shape. The recent discovery of the boson number odd-even effect in 196-204Hg proves the correctness of this theory [26], and also indicates the emergence of previously completely unexpected phenomena in heavy nuclei. These strange color-related phenomena may be very difficult to detect in light nuclei. It is not appropriate to draw conclusions from mere experiences with light nuclei. Thus the influence of the color confinement in the nucleus should be found in the nucleus with more nucleons, especially the medium-heavy nuclei.
@@
How to understand the color confinement in the nucleus is important for the following studies. Equation (1) is the starting point of the discussions. In the simplest case, the representations of the nucleons are (1,0) and (0,1). For more complex situations, (2,0) (0,2) and (1,1) will also need to be considered. The key thing is how to extract the mean field, short-range strong correlations and long-range correlations from all possibilities of the color confinement, the equation (4). This is an interesting and urgent question. It should be noticed that if this division rule exists, it must be sensitive to the nucleon number $A$. We expect that this may have something to do with the magic numbers and the reduction of the SU(3) second-order interaction.
@@
Currently, the main tool for nuclear force researches is the effective field theory [38, 51, 52, 53]. In their discussions, the color degree of freedom is not considered, which means that the quarks and gluons are simply bound to the nucleons. The color confinement in the nucleus is not considered in these studies. It can be seen from this Letter that for a complete understanding of the nuclear force, even at low energy, quarks and gluons need to be considered. How to combine the idea of the effective field with equation (1) is an important work. Just as Miller said:“ all three scales (the nuclear size, the internucleon separation distance and the nucleon size) are must be understood to truly understand the physics of nuclei.” [35] If the division equation (4) exists, the idea of the effective field theory is still applicable, only adding a new limitation.
@@
In conclusion, a simple explanation for the absence of the spherical nuclei is given based on the EMC effect, short-range strong correlations, the Cd puzzle and other new experimental results, and the calculation of the SU3-IBM. It can be obtained that the Cd puzzle is the result of the color confinement in the nucleus. It is clearly pointed out that the color degree of freedom is necessary even in low energy excitation of nuclear structure. The correlations between the low energy and high energy behaviors in the nucleus does exist, and this is the first time to be found here, but what it is requires more further studies.
@@
Here a new theoretical framework to understand nuclear forces is established phenomenologically. Since this explanation is very simple and rigorous, the new connection deserves serious consideration. It does not conflict with previous effective field theories. The possibility of the color degree of freedom between the nucleons is considered here. The SU3-IBM supports this possibility and gives a guide for future theoretical developments. In future, the possible results of this new perspective will be discussed.

### v3

@@ Abstract
Since the EMC effect increases with the number of nucleons, it can be guessed that the effect of this color degree of freedom is more obvious for heavy nuclei. The fact is true. The SU3-IBM explains the prolate-oblate asymmetric phase transition in the Hf-Hg region, where the SU(3) third-order Casimir operator describes the oblate shape. The recent discovery of the boson number odd-even effect in 196-204Hg proves the correctness of this theory Wang , and also indicates the emergence of previously completely unexpected phenomena in heavy nuclei. These strange color-related phenomena may be very difficult to detect in light nuclei. It is not appropriate to draw conclusions from mere experiences with light nuclei. Thus the influence of the color confinement in the nucleus should be found in the nucleus with more nucleons, especially the medium-heavy nuclei.
@@
How to understand the color confinement in the nucleus is important for the following studies. Equation (1) is the starting point of the discussions. The key thing is how to extract the mean field, short-range strong correlations and long-range correlations from all possibilities of the color confinement, the equation (1). This is an interesting and urgent question. It should be noticed that if this division rule exists, it must be sensitive to the nucleon number $A$. We expect that this may have something to do with the magic numbers and the reduction of the SU(3) second-order interaction. Quantum interference among the hidden-color states is a very attractive possibility.
@@
Currently, the main tool for nuclear force researches is the effective field theory Machleidt24 ; Meissner09 ; Machleidt11 ; Kolck20 . In their discussions, the color degree of freedom is not considered, which means that the quarks and gluons are simply bound to the nucleons. The color confinement in the nucleus is not considered in these studies. It can be seen from this Letter that for a complete understanding of the nuclear force, even at low energy, quarks and gluons need to be considered. How to combine the idea of the effective field with equation (1) is an important work. Just as Miller said:“discovery of new phenomena is not well treated by scale separation techniques because new phenomena are often related to discovering a new relevant scale.” “Understanding the EMC effect involves understanding physics at all three length scales (the nuclear size, the internucleon separation distance and the nucleon size).” “all three scales are must be understood to truly understand the physics of nuclei.” Miller20 If the division equation (4) exists, the idea of the effective field theory is still applicable, only adding a new limitation.
What needs to be emphasized is that although the reason for the appearance of the hidden-color states is still unclear, the low-energy collective behavior may be understood by the perturbative quantum chromodynamics Bakker14 ; Ji86 , emphasizing the exchange of the quarks and gluons among the nucleons. This is very important, if so, some of the conclusions given here can be proved. I look forward to further calculations.
VI 6. Conclusion
@@
In conclusion, a simple explanation for the absence of the spherical nuclei is given based on the EMC effect, the short-range strong correlations, the Cd puzzle and other new experimental results, the calculation of the SU3-IBM and the hidden-color state. It can be obtained that the Cd puzzle is the result of the color confinement in the nucleus. If the hidden-color states exist, the spherical nucleus can not appear. It is clearly pointed out that the color degree of freedom is necessary even in low energy excitation of nuclear structure. The correlations between the low-energy and high-energy behaviors in the nucleus does exist, and this is the first time to be found here, but what it is requires more further studies.
@@
In this Letter a new theoretical framework to understand nuclear forces is established phenomenologically and a possible big picture for the whole nuclear physics is drawn. Since this explanation is very simple and rigorous, the new connection deserves serious consideration. It does not conflict with previous effective field theories. The possibility of the color degree of freedom between the nucleons is considered here. The SU3-IBM supports this possibility and gives a guide for future theoretical developments. In future, the possible results of this new perspective will be discussed.